OpenStudy (anonymous):

1-cos(12x)=? A.2cos^2(12x) B.2sin^2(12x) C.2cos^2(6x) D.2sin^2(6x)

OpenStudy (callisto):

1-cos(12x) = 1- (1-2sin^2 (6x)) = 1-1 + 2sin^2 (6x) =?

OpenStudy (anonymous):

b?

OpenStudy (callisto):

Nope :|

OpenStudy (anonymous):

d

OpenStudy (callisto):

Yes

OpenStudy (anonymous):

Thank you

OpenStudy (callisto):

Welcome!~